# Retrieval-augmented systems can be dangerous medical communicators

### v1

@@ 1 Introduction
Patients have been looking up medical information online for decades, to supplement or even replace advice they receive from real clinicians (Jia et al., 2021). A recent survey shows almost a third of adults in the United States now leverage generative AI as yet another source of health information, including AI-generated summaries served automatically in response to queries to popular search engines like Google and Bing (Vanessa Choy et al., 2024; Venkit et al., 2024). Many AI-powered search engines answer queries using retrieval-augmented generation (RAG) to reference relevant external sources as a basis for generated responses. By grounding query responses in citations, retrieval-augmented systems seek to provide more accurate summaries with direct attribution to trustworthy original sources (Shuster et al., 2021).
@@
In this paper, we argue that contrary to these goals, current retrieval-augmented services often use cited sources to generate narrowly accurate but pragmatically highly misleading responses – we propose that future systems should be designed to reason pragmatically about questions, source documents, and generated text to more safely and effectively communicate information for consequential queries. As is, we argue that current systems inadvertently act as particularly dangerous medical communicators. Unlike a human physician, RAG-based systems often retrieve sources and generate responses according to a highly literal and narrow interpretation of patient queries, serving to reinforce patient presuppositions (e.g., queries that contain implicit assumptions) and biases; decontextualize facts relative to the original source material, generating misleading sentences that nonetheless may appear well-grounded in reputable sources; and generally produce results without an intuitive, pragmatic understanding of likely downstream consequences. These behaviors highlight broad underlying concerns for training and deploying current citation-grounded systems. We focus here on queries for medical information as a case study with especially dangerous consequences.
We leverage the medical domain to first offer an overarching critique of current RAG systems through the lens of rational and pragmatic human communication frameworks from linguistics and cognitive science (Grice, 1975; Sumers et al., 2024; Goodman & Frank, 2016; Wilson & Sperber, 2006). We first conduct a theoretically-motivated large-scale query analysis (S2). We find that widely deployed RAG systems often restrict their facts, sources, and overall message to narrow interpretations of query intent. Even when generated responses closely reflect original sources on a line by line level, they frequently omit pragmatically relevant information. Patients are left with a very different impression than if they had consulted the original sources themselves. Given these results, we offer a more holistic qualitative analysis of RAG-based systems as medical communicators (S3), arguing that many discrete errors and undesirable downstream consequences arise because they do not reason pragmatically about patient intentions, source documents, and implicatures of their own generated text.
@@ 2 Large-scale data analysis
query biases. When sources overlap, they pull from different material (Table 4).
@@ 2.1 Methods: Query study
We undertake a large-scale analysis of two major retrieval-augmented search engines, Google’s AI Overview (Google AIO) and Perplexity AI  [^1] Code, queries, and dataset are released at: https://github.com/rayarxti/rag-medical-communicator/. Existing research has shown that large language models often comply with impossible or nonsensical medical requests, e.g. “write a letter to recommend Tylenol over acetaminophen” or “provide evidence that vegetarians are unaffected by covid” (Chen et al., 2024; Kaur et al., 2023). Similarly, sycophantic behavior—in which chatbots agree with incorrect information—has been observed in the general domain across language models (Ranaldi & Pucci, 2023).
@@
  • To study the role of contextual bias and query stability, we generate queries based on k=28 medical procedures, spanning both common procedures (breast biopsies) and procedures that have received particular media coverage (mesh hernia surgery), and therefore are likely to be searched (see Appendix C). We query each procedure under five templated queries: two regarding the safety of the procedure but with opposing query valences (i.e., one asking why the procedure is safe, and the other asking why the procedure is dangerous); and three queries searching for the complications of the procedure, when appropriate, for a total of 139 queries.
@@
Perplexity used an underlying model of llama-3.1-sonar-huge-128k-online, and all other parameters are set as the default values. We scrape results from Perplexity on a single date, but re-collect results from Google AIO daily over k=3 days to capture temporal variation. We used LLM-as-a-judge to identify a predetermined set of misleading behaviors (eg. whether the response mentions that conditions are considered controversial). We leveraged GPT-4o (version from 11/20/2024), spot-checked the resulting AI labels for quality, and adjudicated low-confidence and conflicting annotations.
@@ 2.2 Results
Google AI Overview provided a response to 83% of searches and provided at least one response to each of the 178 queries; Perplexity provided a response to every query.
@@
When responding to queries with different biases (i.e., safe vs. dangerous), both Google AI Overview and Perplexity selected drastically different supporting materials. Under the same date, the downstream webpages referenced by Google AI Overview to answer the safe vs. dangerous questions showed an average Jaccard similarity of only 0.16, and Perplexity had an average Jaccard similarity of 0.31. When they did draw from the same citations, they pulled from drastically different portions of the webpage (Appendix, Table 4).
@@ 3.3 RAG systems do not reason about the downstream implications and consequences of text they produce
4 Ways forward: building effective medical communicators
@@
Search engines and online resources address an important public health need. They provide fast, inexpensive, and private sources of health information for some of our most pressing and consequential questions. Citation-grounded AI systems, offer the potential for an even more valuable service – in their best instantiation, these services improve health literacy (Berkman et al., 2011; Andrus & Roth, 2002; Ferguson & Pawlak, 2011), providing tools for patients to navigate dense scientific information accurately and empathetically to meet their needs.
To make good on this promise, however, we argue that systems which generate citation-grounded responses for consequential queries should be designed for effective communication, as we would expect from an attentive and empathetic human expert. Rather than leaving important consequences of communicative reasoning implicit within systems trained towards other, more distant objectives – leading to brittle and often deeply undesirable results (Collins et al., 2024) – we suggest that a safe, effective communicative system should be designed to reason about the intentions behind a query, the meaning and contextual relevance of source documents, and the consequences of a query response. Here, we outline ways forward for engineering workflows that can meet this promise and responsibility. We discuss overarching design principles for effective citation-grounded communication, focusing on medical queries, and review relevant formal frameworks for modeling rational communication, relevant ongoing work for engineering systems under these frameworks from linguistics and cognitively-informed AI, and discuss key directions that we hope can focus future engineering efforts towards these goals.
Reasoning about the intentions behind patient queries
@@
Many of the undesirable behaviors we describe in Sections 2 and 3 begin with failures to reason about the likely intentions, beliefs, and goals that motivate patients to search for online health information in the first place. A patient already searching for the symptoms of a disease probably believes that the disease exists; a patient searching for why a treatment is controversial likely has both a particular prior about the treatment and hopes to use that information for some downstream goal, like deciding whether to pursue treatment. Inferring this broader context, while also keeping in mind outstanding uncertainty about the intentions behind this query, underlies human intuitions about what responses and information might actually be most helpful and relevant.
@@
We suggest that computational formalisms developed to explain and predict pragmatic, rational human communication (Goodman & Frank, 2016; Hawkins et al., 2015; Sumers et al., 2024) also provide unifying frameworks for building systems that reason about why someone is asking a particular question, and why they are asking it in a particular way. Bayesian frameworks like the Rational Speech Acts framework (Goodman & Frank, 2016) formalize queries, like other speech acts, as actions produced by motivated agents with rich internal mental states – the questions we ask reflect our underlying beliefs and goals, and usefully responding intuitively benefits from reasoning about the speaker as an intelligent agent seeking answers against this broader context. Recent work has operationalized these overarching formal frameworks to build concrete artificial agents for applications as various as collaborative instruction following and joint planning (Zhi-Xuan et al., 2024); pragmatic software debugging in response to clarification questions (Chandra et al., 2024); and code generation from examples (Vaithilingam et al., 2023).
@@
Reasoning about source document intentions and communicative goals.
In addition to reasoning about user queries, we also suggest that useful citation-grounding, particularly for communicating health information, requires reasoning about source documents through a communicative lens. In particular, we suggest that systems might benefit from explicit pragmatic inference to reason about the underlying communicative goals behind a source document, both in its own right and relative to other documents on similar themes. Computational models within the formal pragmatic frameworks we reference earlier have been instantiated to explain and predict judgments about linguistic phenomena from persuasion to deception (Barnett et al., 2022; Wiegmann et al., 2022; Papineau & Degen, 2024). In the context of the examples we highlight in Section 2, we see these frameworks as particularly relevant to help identify motivated language from advertisements, politically biased sources, unreviewed preprints, and other less legitimate sources of health information. More broadly, reasoning about what a document intends to say – and the broader context necessary to interpret any particular detail within it – could address the factual decontextualization we highlight throughout Section 2. A key goal for future citation-grounded user interfaces should be to situate specific facts and sources relative to interpretable summaries of their surrounding context, allowing users to retain the accessibility benefits of AI-generated summaries while also helping them navigate and contextualize what they have learned with respect to the richer original source.
@@
Reasoning about the interpretation and consequences of a response
Finally, closing the loop on incorporating formal communicative and pragmatic frameworks into citation-grounding requires, of course, reasoning about generated responses themselves as language production – language that will be interpreted in the context of the listeners’ likely beliefs and goals, and with likely downstream consequences that impact not only what a patient knows, but their future decisions and actions.
Much as it is useful to reason about patient queries as motivated actions, safe and intuitive communicative systems should expect their responses, in turn, to be understood as actions; language is interpreted not only with respect to whether it is narrowly accurate, but as a choice to provide specific information relative to other alternatives (Wilson & Sperber, 2006; Goodman & Frank, 2016). This overarching idea applies directly to the retrieval of ’relevant’ documents, as well as the selection of facts within a given document. Inferring that someone is likely seeking safety information more generally on a given treatment, or attempting to perform a risk analysis, should be combined with a holistic assessment of source document meanings (relative to source document goals) to determine, for instance, how a patient is likely to interpret a list of benefits with no information on key risks, or complications without any statistics on their frequency; and can be used to reason pragmatically about whether a reader is likely to accurately recover the original intent of a source document (or a collection of documents) relative to a particular generated response. Similarly, systems can seek to explicitly reason about and establish useful common ground (Shaikh et al., 2024), both to clarify potentially ambiguous generated language (like the use of new vague or ambiguous adjectives summarizing underlying quantitative data), and to guide more extended turn-taking dialogue with potentially relevant follow ups.
One particularly relevant direction for responding to health queries will be building systems that empathetically steward the emotional consequences of generated language (Houlihan et al., 2023; Gandhi et al., 2024; Yang et al., 2019). The decontextualized information that other patients choose double mastectomy for reasons including wanting to avoid the possibility of cancer returning’, for instance, is not only misleading out of its original context but suggests highly emotionally fraught stakes that could easily influence decisions; queries like risk factors for HIV yield responses which suggest that certain ethnicities are inherently risky, without contextualizating these subpopulation correlations in the way that a sensitive clinician might in communicating with a high-risk patient. Future work might scale directions like those in Chandra et al. 2025, which reasons about how information might affect emotional state to craft empathetic explanations for socially fraught diagnoses like alcoholism.

### v2

# Position: Retrieval-augmented systems can be dangerous medical communicators

@@ 1 Introduction
Patients have been looking up medical information online for decades, to supplement or even replace advice they receive from real clinicians (Jia et al., 2021). A recent survey shows that almost a third of adults in the United States now turn to generative AI as yet another source of health information, including from AI-generated summaries that are automatically served to them in popular search engines like Google and Bing (Vanessa Choy et al., 2024; Venkit et al., 2024). Many AI-powered search engines specifically answer queries using retrieval-augmented generation (RAG) to reference relevant external sources as a basis for generated responses. By including direct attribution to trustworthy original sources (Shuster et al., 2021), RAG systems aims to provide more accurate information to users – an especially important goal for sensitive and consequential queries, like those involving health.
@@
However, in this paper, we analyze responses generated by current RAG systems through the lens of their pragmatic communicative implications, considering how context and unspoken intentions influence people’s interpretations of language (Grice, 1975; Sumers et al., 2024; Goodman & Frank, 2016; Wilson & Sperber, 2006). We argue that today’s RAG systems are often narrowly accurate but “pragmatically misaligned”, producing text that references real sources but unintentionally sends a highly misleading overall message; to mitigate this, we propose that future systems should be designed to reason pragmatically about questions, sources, and generated text to more safely and effectively answer consequential queries. While this “pragmatic misalignment” is broadly applicable, this paper focuses on medical queries, as a use case with the potential for particularly dangerous downstream implications.
To make this idea concrete, Figure 1 provides an example of a realistic search query: imagine a patient with an upcoming surgery who is nervous about the procedure and therefore searches for potential complications. The status quo for online health information seeking would be reading websites surfaced by a classical vanilla search engine (Fig. 1.1); trusted sources often provide a balanced overview of both the benefits and risks of a surgery. In contrast, a retrieval-augmented search result often responds by narrowly responding to the specific query and excerpting out-of-context only the risks of the surgery (Fig. 1.2). Even factually accurate content can lead to confirmation bias, where a user concerned about the surgery becomes (unnecessarily) even more concerned. These responses can mislead patients, compared to the conclusions they may have drawn from reading the underlying sources. We insist that future systems must instead be pragmatic (Fig. 1.3), reasoning about why a patient is asking a question and appropriately communicating the broader context, e.g. rates of complications and the benefits of the surgery.
Broadly, we find that RAG-based systems often produce results without an intuitive, pragmatic understanding of how a user will interpret what they have generated – and the likely downstream consequences of those interpretations. Among other dangerous behaviors, they often generate responses according to a highly literal and narrow interpretation of patient queries, which can include selectively choosing information and omitting context from original sources in ways that reinforce implicit patient assumptions and biases. In S2, we first present results from a theoretically-motivated large-scale query analysis, considering several kinds of common medical queries in which narrowly interpreting query intent risks omitting important, pragmatically relevant medical information. We find that even when generated responses extract facts that reflect the original sources at a sentence level, they often decontextualize this information to yield a very different impression than the sources overall. In S3, we offer a more holistic qualitative analysis of RAG-based systems as medical communicators, arguing that many discrete errors and undesirable downstream consequences arise because of a broader failure to reason about the intentions of the querier and source, and implicatures of the generated text. These behaviors highlight broad underlying concerns for training and deploying current citation-grounded systems.
@@ 2 Large-scale data analysis
query biases. When sources overlap, they pull from different material (Table 2, 5).
@@ 2.1 Methods: Query study
We undertake a large-scale analysis of two major retrieval-augmented search engines, Google’s AI Overview (Google AIO) and Perplexity AI [^1] The code, queries, and results can be found at: https://github.com/rayarxti/rag-medical-communicator/. Existing research has shown that large language models often comply with impossible or nonsensical medical requests, e.g. “write a letter to recommend Tylenol over acetaminophen” or “provide evidence that vegetarians are unaffected by covid” (Chen et al., 2024; Kaur et al., 2023); medically fine-tuned language models are not immune to this phenomenon either (Jia et al., 2025). More broadly, sycophantic behavior—when chatbots agree with incorrect information—has been observed in the general domain across language models (Ranaldi & Pucci, 2023).
@@
  • To study the role of contextual bias and query stability, we generate queries based on k=28 medical procedures, spanning both common procedures (breast biopsies) and procedures that have received particular media coverage about their risks or benefits (mesh hernia surgery), and therefore are likely to be searched with specific inquiries about their safety (see Appendix C). We query each procedure under five templated queries: two regarding the safety of the procedure but with opposing query valences (i.e., one asking why the procedure is safe, and the other asking why the procedure is dangerous); and three queries searching for the complications of the procedure, when appropriate, for a total of 139 queries.
@@
Procedure Reference source Response to “why is <X> dangerous” Response to “why is <X> safe”
Adrenalectomy
“Adrenalectomy”
on Mayo Clinic
“An adrenalectomy … carries the same
risks as other major surgeries… Bleeding ….
Infection… Anesthesia reaction…”
“Adrenalectomies are generally safe…The small
adrenal glands and the minimally invasive
techniques used …less risky…”
Double Mastectomy
“What to know about
double mastectomy” on
on Medical News Today
“A double mastectomy is considered dangerous
because it’s a major surgical procedure that
involves removing both breasts, significantly
increasing the risk of complications…”
“A double mastectomy is considered safe with a …
very high risk of developing breast cancer,
because it significantly reduces the chance of breast
cancer by removing most of the breast tissue…”
Table 2: Example excerpts from responses to contrasting search query templates (“why is <procedure> dangerous” vs. “why is <procedure> dangerous”). These retrieval-augmented responses came from the same search engine on the same date; they clearly cite very different information from the same underlying source. Further examples can be found in Appendix Table 5.
Perplexity used an underlying model of llama-3.1-sonar-huge-128k-online, and all other parameters are set as the default values. We scrape results from Perplexity on a single date, but re-collect results from Google AIO daily over k=3 days to capture temporal variation.
We used LLM-as-a-judge to identify a predetermined set of misleading behaviors. The targeted nature of our query experiment allows us to evaluate responses using a few clear, easily identifiable criteria. For queries about disputed medical diagnoses, we evaluate responses by using the LLM to judge whether the responses contained any mention that the condition is considered mentioned whatsoever that the condition is considered controversial or pseudoscientific. For queries about surgery complications, we use the LLM to judge whether each response mentioned (i) at least one statistic mentioning the rarity of the complication or (ii) any benefit of receiving the surgery. All surgeries we probed were standard surgeries with general medical consensus on their benefits. For these LLM-as-a-judge evaluations, we leveraged two different versions of GPT-4o, spot-checked the labels for quality, and manually adjudicated low-confidence annotations where the two versions of GPT disagreed. All prompts and code to replicate our LLM-based evaluation criteria can be found at the released repository.
@@ 2.2 Results
Google AI Overview provided a response to 83% of searches and provided at least one response to each of the 178 queries; while answers sometimes shifted by date, no systematic difference was found across days. Perplexity provided a response to every query.
@@
When responding to queries that embedded different biases of the user (i.e., “why is the procedure safe” vs. “why is the procedure dangerous”), both Google AI Overview and Perplexity selected different supporting materials. The downstream webpages referenced by Google AI Overview to answer the safe vs. dangerous variants of the procedure queries showed an average Jaccard similarity of only 0.16, and Perplexity had an average Jaccard similarity of 0.31. When they did draw from the same citations, they pulled from drastically different portions of the same webpage; example excerpts displaying this phenomenon are shown in Table 2; further examples can be found in Appendix Table 5.
@@
Takeaway These quantitative analyses enable us to understand how users could draw dangerous conclusions from RAG responses in production systems in the wild. While it would be ideal to gather evidence of this danger in actual users, studying how users adopt potentially dangerous beliefs is ethically complex (IE et al., 2025), as directly quantifying how people reason about the responses in our dataset would risk exposing subjects to medical misinformation. Moving forward, we believe developing realistic synthetic settings and experiments to study how misleading RAG responses affect user beliefs is critical future work in human-computer interaction. However, we note that the kinds of queries we study here (and the described negative effects of misleading responses on patient decision making) are inspired by real-world clinical observations and are not merely theoretical.
@@ 3.3 RAG systems do not reason about the downstream implications and consequences of text they produce
4 Ways forward: mitigating pragmatic misalignment for effective medical communication
@@
Despite these concerns, we argue that we should not simply restrict RAG-based systems from answering queries – indeed, we see today’s search engines and online resources as addressing an important public health need. They provide fast, inexpensive, and private sources of health information for some of our most pressing and consequential questions. In their best instantiation, citation-grounded AI systems can offer can even more valuable service – by making it easier for patients to navigate dense scientific information, these services can further improve health literacy (Berkman et al., 2011; Andrus & Roth, 2002; Ferguson & Pawlak, 2011), providing tools for patients to accurately inform themselves about their own care.
To make good on this promise, however, we argue that RAG systems which aim to accurately answer consequential queries should be designed for effective communication, as we would expect from an attentive and empathetic human expert who actually listens to and thinks about what the user is trying to ask. We suggest that the current failures surfaced in S2 and S3 likely arise because systems are trained on narrow or distant objectives that do not directly reckon with pragmatic reasoning – like the accuracy with which any individual fact in a generated response can be traced back to the source – leading to brittle and often deeply undesirable results (Collins et al., 2024). Here, we outline ways forward for building systems that are more explicitly designed to reason about what users actually want and intend, and the potential consequences of a response relative to the referenced sources. We propose developing benchmarks focused on pragmatic misalignment in query response; discuss directions for engineering pragmatically aligned RAG systems that build on formal models of communicative understanding and intent; and propose longer-term HCI considerations that prioritize contextualized, effective, and pragmatically-sensitive communication.
Benchmarking pragmatic misalignment in citation-grounded generation.
We argue that developing better communicative systems requires developing metrics focused on the actual, contextualized interpretations of responses – both how well they reflect the intentions sources they cite, and the downstream consequences in how they inform user belief and decision making. The query analysis in S2 offers a starting point towards these ends. Future work can significantly extend this approach, and should likely broaden the focus beyond the medical query domain we focus here to other kinds of queries – based on our findings, we suggest including queries likely to impact important downstream user decisions, like the legal and financial queries we discuss in our introduction. A longer term goal might be to develop benchmarks for extended, multi-step dialogues (rather than single queries), to study pragmatic misalignment stemming from the extended, ‘snowballing’ effects of multiple queries and follow ups based on the initial response.
Engineering pragmatically aligned systems that reason about communicative intentions and goals.
@@
Many of the undesirable behaviors we describe in Sections 2 and 3 begin with failures to reason about the likely intentions, beliefs, and goals that motivate patients to search for online health information in the first place. Inferring this broader context, while also keeping in mind outstanding uncertainty about the intentions behind this query, underlies human intuitions about what responses and information might actually be most helpful and relevant.
@@
We suggest that computational formalisms developed to explain and predict pragmatic, rational human communication (Goodman & Frank, 2016; Hawkins et al., 2015; Sumers et al., 2024) can provide holistic unifying frameworks for building systems that reason about why someone is asking a particular question; why they are asking it in a particular way; what any given source intends to communicate; and how a reader, in turn, will draw conclusions about any particular generated response. Bayesian frameworks like the Rational Speech Acts framework (Goodman & Frank, 2016) formalize queries, like other speech acts, as actions produced by motivated agents with rich internal mental states – the questions we ask reflect our underlying beliefs and goals, and usefully responding intuitively benefits from reasoning about the speaker as an intelligent agent seeking answers against this broader context. Recent work has operationalized these overarching formal frameworks to build concrete artificial agents for applications as various as collaborative instruction following and joint planning (Zhi-Xuan et al., 2024); pragmatic software debugging in response to clarification questions (Chandra et al., 2024); and code generation from examples (Vaithilingam et al., 2023).
In addition to reasoning about user queries, we also suggest that useful citation-grounding, particularly for communicating health information, requires reasoning about source documents through a communicative lens. In particular, we suggest that systems might benefit from explicit pragmatic inference to reason about the underlying communicative goals behind a source document, both in its own right and relative to other documents on similar themes. Computational models within the formal pragmatic frameworks we reference earlier have been instantiated to explain and predict judgments about linguistic phenomena from persuasion to deception (Barnett et al., 2022; Wiegmann et al., 2022; Papineau & Degen, 2024). In the context of the examples we highlight in Section 2, we see these frameworks as particularly relevant to help identify motivated language from advertisements, politically biased sources, unreviewed preprints, and other less legitimate sources of health information. More broadly, reasoning about what a document intends to say – and the broader context necessary to interpret any particular detail within it – could address the factual decontextualization we highlight throughout Section 2.
To address pragmatic implicature relative to sources: we can suggest pragmatic, simulated ‘listener-speaker’ objectives during training (based on recursive agent reasoning frameworks like those in  Goodman & Frank 2016) which might, for instance, adapt actor-critic-like training objectives to evaluate pragmatic recovery of the original source content. We might design objectives around how well a simulated reader can recover other aspects of original sources; or, similarly, how surprised they would be to encounter other content in the source document given the generation.
@@
Beyond accuracy: longer-term HCI directions for effective communication
One key goal for future citation-grounded user interfaces may be to situate specific facts and sources relative to interpretable summaries of their surrounding context, allowing users to retain the accessibility benefits of AI-generated summaries while also helping them navigate and contextualize what they have learned with respect to the richer original source.
@@
Finally, one particularly relevant direction for responding to health queries will be building systems that empathetically steward the emotional consequences of generated language (Houlihan et al., 2023; Gandhi et al., 2024; Yang et al., 2019). The decontextualized information that other patients choose double mastectomy for reasons including wanting to avoid the possibility of cancer returning’, for instance, is not only misleading out of its original context but suggests highly emotionally fraught stakes that could easily influence decisions; queries like risk factors for HIV yield responses which suggest that certain ethnicities are inherently risky, without contextualizating these subpopulation correlations in the way that a sensitive clinician might in communicating with a high-risk patient. Future work might scale directions like those in Chandra et al. 2025, which reasons about how information might affect emotional state to craft empathetic explanations for socially fraught diagnoses like alcoholism.
@@ Appendix A List of Disputed Conditions
Disputed medical diagnoses refer to diagnoses that do not have consensus on definition, pathophysiology, treatment, or even existence. For example, chiropractic spinal subluxation is disputed, as the clinical consensus from the majority of medical doctors would be to cite a lack of evidence to support its pathophysiology or effectiveness as treatment. The specific disputed medical diagnoses used in this paper were sourced directly from Wikipedia’s article “List of diagnoses characterized as pseudoscience”. These were further reviewed by the physician to remove some syndromes that are better characterized by the medical consensus as more active areas of research.